Please show work if possible and answer both examples for a thumbs up. Thank you!

A stock just paid a dividend of $2.12. The dividend is expected to grow at 20.87% for three years and then grow at 3.00% thereafter. The required return on the stock is 10.91%. What is the value of the stock?

A stock just paid a dividend of $2.33. The dividend is expected to grow at 20.09% for three years and then grow at 4.19% thereafter. The required return on the stock is 12.12%. What is the value of the stock?

Answer format: Currency: Round to: 2 decimal place
Answer #1

Part 1:

Dividend just paid =2.12
Dividend in year 1=2.12*(1+Growth)=2.12*(1+20.87%)=2.562444
Dividend in year 2=2.12*(1+20.87%)*(1+20.87%)=3.097226063
Dividend in year 3=2.12*(1+20.87%)*(1+20.87%)*(1+20.87%)=3.743617142


Price beyond year 3 is equal to [(Dividend in year 3)*(1+Constant growth)]/(Required return-constant growth)
=3.743617142*(1+3%)/(10.91%-3%)
=3.855925656/0.0791=48.74747985

Current price=(Dividend in year 1)/(1+Required return)^1+(Dividend in year 2)/(1+Required return)^2+(Dividend in year 3)/(1+Required return)^3+(Price beyond year 3)/(1+Required return)^3

=(2.562444)/(1+10.91%)^1+(3.097226063)/(1+10.91%)^2+(3.743617142)/(1+10.91%)^3+(48.74747985)/(1+10.91%)^3

=2.31038139+2.517859514+2.7439697+35.7305789
Current value of the stock=43.3027895 or $43.30 (Rounded to two decimal places)

Part 2:
Dividend just paid =2.33
Dividend in year 1=2.12*(1+Growth)=2.12*(1+20.09%)=2.545908

Dividend in year 2=2.12*(1+20.09%)*(1+20.09%)= 3.057380917

Dividend in year 3=2.12*(1+20.09%)*(1+20.09%)*(1+20.09%)= 3.671608743
Price beyond year 3 is equal to [(Dividend in year 3)*(1+Constant growth)]/(Required return-constant growth)
=3.671608743*(1+4.19%)/(12.12%-4.19%)
=48.24021626

Current price=(Dividend in year 1)/(1+Required return)^1+(Dividend in year 2)/(1+Required return)^2+(Dividend in year 3)/(1+Required return)^3+(Price beyond year 3)/(1+Required return)^3
Current price=(2.545908)/(1+12.12%)^1+(3.057380917)/(1+12.12%)^2+(3.671608743)/(1+12.12%)^3+(48.24021626)/(1+12.12%)^3

=(2.545908)/(1+12.12%)^1+(3.057380917)/(1+12.12%)^2+(3.671608743)/(1+12.12%)^3+(48.24021626)/(1+12.12%)^3
=2.270699251+2.43211089+2.604996404+34.22630206
Current value of the stock=41.53410861 or $41.53 (Rounded to two decimal places)